OpenStudy (anonymous):

a person places rm20,000 in a saving account which pays 5 percent interest per annum , compounded continuously. a)derive a differential equation that model the initial-value problem b)find the solution of the differential equation c)find the amount in the account after three years d)find the time required for the account to double in value,presuming no withdrawal and no additional deposits

OpenStudy (paxpolaris):

\[\huge A=P \cdot e^\left( rt \right)\] P=20,000 r= 5/100 = 0.05

OpenStudy (anonymous):

thank you PaxPolaris :)